If sec theta + tan theta = x . Then find the value of sec theta
 

Grade:10

1 Answers

Arun
5 years ago
As we now
 
sec^2 A – tan^2 A = 1
 
(secA + tanA) (secA – tanA) = 1
 
secA + tanA = 1 / (secA – tanA) = 1/ x
 
hence
 
add both the equations
 
2 sec A = x + 1/ x
 
secA = x^2 + 1 / 2x